# Gibbs sampling, exact sampling, and distribution function evaluation for the generalized inverse Gaussian distribution

### v2

@@ 2 Data-augmented Gibbs sampler
If $p>-1/2$, then
$f_{X\mid Y}(x\mid y)$ $=\mathrm{InvGauss}(x\mid\sqrt{(b+2y)/a},b+2y)$
By Proposition 1, we can define a joint PDF $f_{X,Y}(x,y)=f_{X\mid Y}(x)f_{Y}(y)$ whose marginal $f_{X}(x)$ is $\mathrm{GIG}(x\mid p,a,b)$. We have established that $X\mid Y=y$ is inverse Gaussian. The Proposition below states that $Y\mid X=x$ is gamma-distributed.
@@
We can construct a data-augmented Gibbs sampler by sampling iteratively from the conditional distributions $X\mid Y=y$ and $Y\mid X=x$ (see Algorithm 1). Only the draws from $X\mid Y=y$ are relevant if we are interested in sampling from the GIG$(p,a,b)$. The case $p=-1/2$ is not included in Propositions 1 and 2, but the $\mathrm{GIG}(-1/2,a,b)$ is simply $\mathrm{InvGauss}(\sqrt{b/a},b)$ [22].
There are efficient algorithms to generate random draws from the inverse Gaussian and gamma distributions. To generate inverse Gaussian draws, one can use the algorithm described in [28], whereas for the gamma distribution, one can use standard algorithms such as those proposed in [2].
@@
The GIG$(p,a,b)$ often arises as a full conditional distribution in Gibbs sampling algorithms. There are countless examples from the literature on global-local shrinkage priors [3, 33, 7, 16], graphical models [23, 29], stochastic volatility models [6], Bayesian nonparametrics [13], smoothing splines [27], and other areas. The Gibbs sampler described in Algorithm 1 can be embedded within an encompassing MCMC algorithm when simulation from a GIG$(p,a,b)$ is required. The following example offers a simple stylized illustration.
Example 1.
Suppose that we observe independent and identically distributed data $y_{i}\mid\mu,\sigma^{2}\sim\mathrm{Normal}\left(\mu,\sigma^{2}\right)$, $i\in\{1,2,\,...\,,n\}$, with $\mu$ and $\sigma^{2}$ unknown. We assign independent priors to $\mu$ and $\sigma^{2}$ with $\mu\sim\mathrm{Normal}\left(\theta_{0},\tau_{0}^{2}\right)$ and $\sigma^{2}\sim\mathrm{GIG}\left(p_{0},a_{0},b_{0}\right).$ Letting $\mathbf{y}=\left(y_{1},\ldots,y_{n}\right)^{\top}$ and $\bar{y}=\frac{1}{n}\sum_{i=1}^{n}y_{i},$ we can simulate from the posterior distribution of $\left(\mu,\sigma^{2}\right)$ by iterating between the full conditional distributions
where
and
If $p_{n}>-1/2,$ the full conditional distributions are
The Gibbs sampler in Algorithm 1 is geometrically ergodic, which is a property that quantifies the rate of convergence of the Gibbs sampler to its stationary distribution. It also ensures that inferences made with draws from the chain are well-behaved. In particular, if $\{X_{i}\}_{i=1}^{n}$ are draws from the chain and we want to estimate a function $f(X)$ with finite $\mathbb{E}[f(X)^{2+\varepsilon}]$ for some $\varepsilon>0$, geometric ergodicity ensures that $\sum_{i=1}^{n}f(X_{i})/n$ is approximately normal [10].
More formally, geometric ergodicity can be defined as follows (our exposition mirrors that in [21]). Let the $k$-step ahead transition kernel of the Markov chain be
$P^{k}[(x,y),A]=P[(X_{i+k},Y_{i+k})\in A\mid(X_{i},Y_{i})=(x,y)],$
where $A$ is a $P^{k}$-measurable set. In our context, $A\in\mathcal{B}(\mathbb{R}^{2}_{+})$, where (+$\mathcal{B}(\mathbb{R}^{2}_{+})$ are the Borel sets on +$\mathbb{R}^{2}_{+}$. Let $\pi(\cdot)$ be the probability measure of the stationary distribution of the chain. Then, geometric ergodicity requires the existence of a function $M(\cdot,\cdot)$ and a constant $t\in(0,1)$ so that
$\sup_{A\in\mathcal{B}(\mathbb{R}^{2}_{+})}|P^{k}[(x,y),A]-\pi(A)|\leq t^{k}M(x,y)$
for all $(x,y)$ in the support of $(X,Y)$, which in our case is +$\mathbb{R}^{2}_{+}$. In other words, the total variation distance between the $k$-step ahead kernel and the stationary distribution decreases exponentially fast in $k$ for all $(x,y)\in\mathbb{R}^{2}_{+}$.
We use Theorem 3.5 in [21] to establish that the Gibbs sampler described in Algorithm 1 is geometrically ergodic. Since the densities involved are all continuous, the proof amounts to showing that there exist functions $f,g:[0,\infty)\rightarrow[1,\infty)$ and constants $j,k,m,n>0$ with $jm<1$ so that
with $C_{d}=\{y:g(y)\leq d\}$ compact for all $d>0$.
Proposition 3.
Algorithm 1 describes a composition Gibbs sampler because $X$ and $Y$ are always sampled in the same order. Random sequence scan and random scan Gibbs samplers can be defined in an analogous manner. In random sequence scan and random scan Gibbs samplers, the order in which $X$ and $Y$ are sampled is random and can change from iteration to iteration (see, for example, [21] for further details). The results in [21] imply that the analogous random sequence scan and random scan Gibbs samplers are also geometrically ergodic.
3 Exact sampling and CDF evaluation for half-integer $p$
The main results of this section reveal a relationship between GIGs with different values of $p.$ Proposition 4 leads to the first algorithm in the literature for exact, rejection-free simulation from the GIG$(p,a,b)$ for half-integer $p.$ Proposition 5 leads to a simple algorithm for evaluating the CDF of the GIG$(p,a,b)$ for half-integer $p.$ The results can be established using Theorem 2.4 of [20].
Proposition 4.
$f_{Y}(y)$ $=w\,\mathrm{GIG}(y\mid p-2,a,b)+(1-w)\,\mathrm{GIG}(y\mid p-1,a,b)$
$w$ $={K_{p-2}(\sqrt{ab})}/{K_{p}(\sqrt{ab})}\in(0,1)$
and $E\sim\mathrm{Exp}(a/2).$
The exact sampling algorithm for half-integer $p$ (Algorithm 2) is a direct consequence of Proposition 4. As described, the algorithm generates a single random number. To draw more than one number, one can wrap Algorithm 2 within a for loop.
@@ 3 Exact sampling and CDF evaluation for half-integer $p$
The algorithm for evaluating the CDF of the GIG for half-integer $p$ (Algorithm 3) relies on the following recurrence formula, which is a consequence of Proposition 4.
Proposition 5.
where
and $\Gamma(\cdot,\cdot)$ is the incomplete gamma function [1].
The CDFs $G_{-1/2}$ and $G_{1/2}$ are both known. The former is the CDF of $\mathrm{InvGauss}(\sqrt{b/a},b)$. The latter can be evaluated with the CDF of an inverse Gaussian because if $X\sim\mathrm{GIG}(-p,b,a)$, then $1/X\sim\mathrm{GIG}(p,a,b)$ [22].
Proposition 5 suggests the implementation of a recursive algorithm that repeatedly calls the function to obtain $G_{p-2}$ and $G_{p-1}$ until reaching the base case. However, the bottom-up approach described in Algorithm 3 is substantially faster for moderate to large $p$ because it avoids redundant computations. The algorithm could be optimized further by storing the evaluations of the Bessel functions as they are computed. We do not pursue this in Algorithm 3 to simplify our exposition.
@@
The base cases of the algorithms involve the inverse Gaussian distribution. Generating random draws from the inverse Gaussian and evaluating its CDF is straightforward. Both algorithms require the evaluation of the modified Bessel function of the second kind for half-integer orders of $p$. This can be done using well-known properties (see, for instance, [1]): (1) $K_{-\nu}(z)=K_{\nu}(z)$, (2) $K_{1/2}(z)=\sqrt{\pi/2z}\,\exp\{-z\}$, and (3) the recurrence formula $K_{\nu+1}(z)=2\nu K_{\nu}(z)/z+K_{\nu-1}(z).$
It is not uncommon that statistical methods require sampling from or evaluating the CDF of the $\text{GIG}(p,a,b)$ for half-integer $p.$ [13] introduce a stick-breaking representation for the normalized inverse Gaussian process [26] whose practical value depends upon being able to efficiently simulate from the GIG for half-integer $p.$ In both [7] and [27], GIG$(p,a,b)$ full conditional distributions have half-integer $p$ parameters for the priors considered. [18] present a data augmentation scheme for models that include gamma functions that requires simulating from a sum of independent GIGs for which $p=-3/2.$ [31] describe several financial and actuarial applications in which one needs to evaluate the CDF of the $\text{GIG}(p,a,b)$ for half-integer $p.$
4 Conclusions
This article introduced representations of the generalized inverse Gaussian distribution that lead to simple algorithms for generating random draws and evaluating its CDF. As future work, it would be interesting to study if there are analogous representations for the matrix-variate version of the generalized inverse Gaussian [4, 9, 17], a probability distribution over symmetric positive-definite matrices that includes the Wishart and inverse Wishart as special cases.
Appendix A Proofs of Propositions
Proposition 1.
$f_{X}(x)=\int_{0}^{\infty}f_{X\mid Y}(x\mid y)f_{Y}(y)\,\mathrm{d}y,$
where $f_{X}(x)$ is the PDF of the $\mathrm{GIG}(p,a,b)$, $f_{Y}(y)$ is a PDF with support on $\mathbb{R}_{+}$, and $f_{X\mid Y}(x\mid y)$ is an inverse Gaussian PDF. The parameters of the distributions depend on whether $p<-1/2$ or $p>-1/2$. If $p<-1/2$, then
$f_{X\mid Y}(x\mid y)$ $=\mathrm{InvGauss}(x\mid\sqrt{b/(a+2y)},b)$
If $p>-1/2$, then
$f_{X\mid Y}(x\mid y)$ $=\mathrm{InvGauss}(x\mid\sqrt{(b+2y)/a},b+2y)$
Proof.
$f_{X}(x)\propto x^{p+1/2}\,\mathrm{InvGauss}(x\mid\sqrt{b/a},b).$
The term $x^{p+1/2}$ can be represented as follows
so we can write
$f_{X}(x)\propto\mathrm{InvGauss}(x\mid\sqrt{b/a},b)\,\int_{0}^{\infty}y^{-(p+3/2)}\,\exp\left(-xy\right)\,\mathrm{d}y.$
If we define
$g(x,y)=\mathrm{InvGauss}(x\mid\sqrt{b/a},b)\,y^{-(p+3/2)}\,\exp\left(-xy\right)\mathbbm{1}(x>0,y>0),$
$X\mid Y=y$ $\sim\mathrm{InvGauss}(\sqrt{b/(a+2y)},b)$
$Y\mid X=x$ $\sim\mathrm{Gamma}(-(p+1/2),x).$
$x^{p+1/2}=\frac{\int_{0}^{\infty}y^{p-1/2}\,\exp(-y/x)\,\mathrm{d}y}{\Gamma(p+1/2)}.$
We can rewrite the PDF of the $\mathrm{GIG}(p,a,b)$ as
$f_{X}(x)\propto\mathrm{InvGauss}(x\mid\sqrt{b/a},b)\,\int_{0}^{\infty}y^{p-1/2}\,\exp\left(-y/x\right)\,\mathrm{d}y.$
By Tonelli’s theorem, the joint PDF of $(X,Y)$ is proportional to
$f_{X,Y}(x,y)\propto y^{p-1/2}\exp\left\{-\sqrt{a(b+2y)}\right\}\,\mathrm{InvGauss}(x\mid\sqrt{(b+2y)/a},b+2y).$
The conditional distributions are
$X\mid Y=y$ $\sim\mathrm{InvGauss}(\sqrt{(b+2y)/a},b+2y)$
$Y\mid X=x$ $\sim\mathrm{Gamma}(p+1/2,1/x).$
Lastly, we find $f_{Y}(y)$. Again, we treat the cases $p<-1/2$ and $p>-1/2$ separately. If $p<-1/2$, then:
$f_{Y}(y)^{-1}$ $=\int_{0}^{\infty}\frac{f_{X}(x\mid y)}{f_{Y}(y\mid x)}\,\mathrm{d}x$
$=\int_{0}^{\infty}\frac{\mathrm{InvGauss}(x\mid\sqrt{b/(a+2y)},b)}{\mathrm{Gamma}(y\mid-(p+1/2),x)}\,\mathrm{d}x$
$=\mathcal{Z}_{0}\,y^{p+3/2}\,\exp(b^{1/2}(a+2y)^{1/2}),$

### v3

@@ A.1 Proofs of Propositions
The case $p>-1/2$ can be handled similarly:
$f_{Y}(y)^{-1}$ $=\int_{0}^{\infty}\frac{f_{X}(x\mid y)}{f_{Y}(y\mid x)}\,\mathrm{d}x$
$=\int_{0}^{\infty}\frac{f_{\mathrm{IG}}(x\,;\,\sqrt{(b+2y)/a},b+2y)}{f_{\mathrm{G}}(x\,;\,p+1/2,1/x)}\,\mathrm{d}x$
where
@@
Proof.
$f_{X,Y}(x,y)\propto f_{\mathrm{IG}}(x\,;\,\sqrt{b/a},b)\,y^{-(p+3/2)}\,\exp\left(-xy\right),$
from which it is straightforward to deduce that
$Y\mid X=x$ $\sim\mathrm{Gamma}(-(p+1/2),x).$
If $p>-1/2$, then
$f_{X,Y}(x,y)\propto y^{p-1/2}\exp\left\{-\sqrt{a(b+2y)}\right\}\,f_{\mathrm{IG}}(x\,;\,\sqrt{(b+2y)/a},b+2y),$
which implies that
$Y\mid X=x$ $\sim\mathrm{Gamma}(p+1/2,1/x).$
Proposition 3.
Proof.
We prove this result by checking that the conditions of Theorem 3.5 in Johnson and Burbank (2015) are satisfied. The proof amounts to showing that there exist functions $f,g:[0,\infty)\rightarrow[1,\infty)$ and constants $j,k,m,n>0$ with $jm<1$ so that
Case $p<-1/2$: Define
$\theta$ $=3/(2b)$
$\mu$ $=3(1/2-p)/(2b)$
$f(x)$ $=1+\alpha(1/x-\theta)^{2},\qquad 0<\alpha<b^{2}/(p^{2}-1/4)$
$\mathrm{Var}(1/X\mid Y=y)$ $={(a+2y)^{1/2}}{b^{-3/2}}+{2}{b^{-2}}$
Rearranging terms, we obtain
$=1+9\alpha/(4b^{2})+\alpha a/b+2\alpha\mu/b+2\alpha(y-\mu)/b$
$\leq 1+9\alpha/(4b^{2})+\alpha a/b+2\alpha\mu/b+g(y),$
$\mathbbm{E}[f(X)\mid Y=y]$ $\leq jg(y)+k$
$j$ $=1$
$k$ $=1+9\alpha/(4b^{2})+\alpha a/b+2\alpha\mu/b.$
Both $j$ and $k$ are positive, which is required by Theorem 3.5 in Johnson and Burbank (2015). Now, we find the conditional expectation of $g$ and bound it:
$\leq 1+9\beta(1/2-p)/(4b^{2})+\beta(p^{2}-1/4)f(x)/\alpha.$
We can rewrite the inequality as
$\mathbbm{E}[g(Y)\mid X=x]$ $\leq mf(x)+n$
$m$ $=\beta(p^{2}-1/4)/\alpha$
$n$ $=1+9\beta(1/2-p)/(4b^{2}),$
where $m,n>0$, as required by Theorem 3.5 in Johnson and Burbank (2015). It remains to show that $jm<1$. Given our choices of $\alpha$ and $\beta$,
$10<{\beta(p^{2}-1/4)}/{\alpha}<$ 1
Case $p>-1/2:$ Let $10<\gamma<$ 1 and
$\theta$ $=1/(2a)$
$\mu$ $=(p+1/2)/(2a)$
$f(x)$ $=1+\alpha(x-\theta)^{2},\qquad\alpha>0$
$=1+\alpha/(4a^{2})+\alpha b/a+{2\alpha}\mu/a+{2\alpha}(y-\mu)/a$
$\leq 1+\alpha/(4a^{2})+\alpha b/a+{2\alpha}\mu/a+\gamma g(y).$
$\mathbbm{E}[f(X)\mid Y=y]$ $\leq jg(y)+k$
$j$ $=\gamma$
$k$ $=1+\alpha/(4a^{2})+\alpha b/a+{2\alpha}\mu/a,$
where $j$ and $k$ are positive, as required by Theorem 3.5 in Johnson and Burbank (2015). Finally, we bound the conditional expectation of $g$:
$=1+\beta\left\{(p+1/2)x^{2}+[(p+1/2)x-\mu]^{2}\right\}$
$=1+\beta(p+1/2)x^{2}+\beta(p+1/2)^{2}(x-\theta)^{2}$
$=\beta(p+1/2)x^{2}+[\beta(p+1/2)^{2}-\alpha](x-\theta)^{2}+f(x)$
$=Q(x)+f(x)$
$\leq c_{0}+f(x),$
$\mathbbm{E}[g(Y)\mid X=x]$ $\leq mf(x)+n$
$m$ $=1$
$n$ $=c_{0},$
Proposition 4.
$f_{Y}(y)$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}w\,f_{\mathrm{GIG}}(y\,;\,p-2,a,b)+(1-w)\,f_{\mathrm{GIG}}(y\,;\,p-1,a,b)}$
$w$ $={K_{p-2}(\sqrt{ab})}/{K_{p}(\sqrt{ab})}\in(0,1)$
and $E\sim\mathrm{Exp}(a/2)$, where $f_{\mathrm{GIG}}(y\,;\,p,a,b)$ is the PDF of the GIG$(p,a,b)$.
Proof.
$f_{X}(x)$ $=\int_{0}^{\infty}f_{X\mid Y}(x\mid y)f_{Y}(y)\,\mathrm{d}y$
where
$f_{Y}(y)$ $=w\,f_{\mathrm{GIG}}(y\,;\,p-2,a,b)+(1-w)\,f_{\mathrm{GIG}}(y\,;\,p-1,a,b)$
$w$ $=\frac{K_{p-2}(\sqrt{ab})}{K_{p}(\sqrt{ab})}.$
Let $a>0,b>0,p>1$, and define:
[…]
The ratio
[…]
is monotone non-decreasing in $x>0$ for $p>1$, so we can apply Theorem 2.4 in Jauch et al. (2023) to write $g(x)$ as a mixture of truncated $\mathrm{Exp}(a/2)$ random variables:
$g(x)=\int_{0}^{\infty}\frac{a}{2}\exp\left\{-\frac{a}{2}(x-y)\right\}\mathbbm{1}(x\geq y)\,f_{Y}(y)\,\mathrm{d}y.$
By Theorem 2.4 in Jauch et al. (2023), the mixing density $f_{Y}(y)$ is
[…]
$=w\,f_{\mathrm{GIG}}(y\,;\,p-2,a,b)+(1-w)\,f_{\mathrm{GIG}}(y\,;\,p-1,a,b).$
$f_{\tilde{X}}(x)$ $=\int_{0}^{\infty}f_{E}(x-y)f_{Y}(y)\,\mathrm{d}y$
$=f_{\mathrm{GIG}}(x\,;\,p,a,b),$
as we wanted to show. ∎
Proposition 5.
where
and $\Gamma(\cdot,\cdot)$ is the incomplete gamma function (Abramowitz and Stegun, 1968).
Proof.
Let $G_{p}$ be the CDF of $\mathrm{GIG}(p,a,b)$. Then,
On the one hand,
$\int_{0}^{x}f_{Y}(y)\,\mathrm{d}y=wG_{p-2}(x)+(1-w)G_{p-1}(x),$
$\int_{0}^{x}\exp\left(\frac{ay}{2}\right)f_{Y}(y)\,\mathrm{d}y=I_{p}(x)$
where $\Gamma(\cdot,\cdot)$ is the incomplete gamma function (Abramowitz and Stegun, 1968). ∎
A.2 Algorithms
In this section, we give pseudocode to implement the algorithms. In Algorithm 3 (CDF evaluation for half-integer $p$), we take a bottom-up approach (a for loop). This is more efficient than a recursive function because it avoids redundant computations.
@@ A.2 Algorithms
Appendix B Reparametrization of the GIG
This section discusses a reparametrization of the GIG and restates all the propositions in the main text in this alternative parametrization.
B.1 Reparametrization
Let $f_{X}(x)$ denote the PDF of $X\sim\text{GIG}(p,a,b)$, given by:
$f_{X}(x)=\frac{\left(a/b\right)^{p/2}}{2K_{p}(\sqrt{ab})}x^{p-1}\exp\left(-\frac{ax+b/x}{2}\right),\quad x>0,$
where $K_{p}(\cdot)$ is the modified Bessel function of the second kind. The parameters are $a>0$, $b>0$, and $p\in\mathbb{R}$.
Consider the reparametrization
$\omega=\sqrt{ab},\quad\eta=\sqrt{b/a},\quad p=p.$
Substituting these into the PDF, the density function becomes:
$f_{X}(x)=\frac{1}{2\eta^{p}K_{p}(\omega)}x^{p-1}\exp\left(-\frac{\omega}{2}\left(\frac{x}{\eta}+\frac{\eta}{x}\right)\right),\quad x>0.$
This reparametrization is useful because generating a random draw from $\text{GIG}(p,a,b)$ is equivalent to generating from $\text{GIG}(p,\omega,\eta=1)$ and then rescaling the draws:
  1. 1.
    Compute $\omega=\sqrt{ab}$ and $\eta=\sqrt{b/a}$.
  2. 2.
    Generate a random draw $Z\sim\text{GIG}(p,\omega,\eta=1)$.
  3. 3.
    Scale $Z$ by $\eta$ to obtain $X=\eta Z$.
Proof: Let $Z\sim\text{GIG}(p,\omega,\eta=1)$ and define $X=\eta Z$, where $\eta=\sqrt{b/a}$. Using the change of variables $z=x/\eta$, the Jacobian is $\mathrm{d}z/\mathrm{d}x=1/\eta$. The PDF of $X$ is:
$f_{X}(x)=f_{Z}\left(\frac{x}{\eta}\right)\frac{1}{\eta}.$
Substituting $f_{Z}(z)$ and simplifying:
$f_{X}(x)=\frac{1}{2K_{p}(\omega)}\left(\frac{x}{\eta}\right)^{p-1}\exp\left(-\frac{\omega}{2}\left(\frac{x}{\eta}+\frac{\eta}{x}\right)\right)\frac{1}{\eta}.$
Combine terms:
$f_{X}(x)=\frac{\eta^{-p}}{2K_{p}(\omega)}x^{p-1}\exp\left(-\frac{ax+b/x}{2}\right).$
Since $\eta^{-p}=\left(a/b\right)^{p/2}$, we recover:
$f_{X}(x)=\frac{(a/b)^{p/2}}{2K_{p}(\omega)}x^{p-1}\exp\left(-\frac{ax+b/x}{2}\right).$
Thus, $X\sim\text{GIG}(p,a,b)$.
B.2 Propositions in the new parametrization
This section restates the propositions using $\omega$ and $\eta$. Propositions 4 and 5 involve the weight $w$, which determines how different GIGs are combined. Unfortunately, the notation $w$ clashes with the new parameter $\omega$, so we redefine
$\tau=w=\frac{K_{p-2}(\omega)}{K_{p}(\omega)}.$
Proposition 1: The PDF of $X\sim\text{GIG}(p,\omega,\eta)$ can be expressed as a mixture of inverse Gaussian distributions:
$f_{X}(x)=\int_{0}^{\infty}f_{X|Y}(x\mid y)f_{Y}(y)\,\mathrm{d}y,$
where:
  • For $p<-1/2$:
    $f_{X|Y}(x\mid y)=f_{\mathrm{IG}}(x\,;\,\sqrt{\omega\eta/(\omega+2\eta y)},\omega\eta),$
    $f_{Y}(y)=\frac{1}{\mathcal{Z}_{0}}\,y^{-(p+3/2)}\exp\left\{-\,\omega\sqrt{1+2\eta y/\omega}\right\},$
    with
    $\mathcal{Z}_{0}=\frac{\sqrt{2}\,\Gamma\bigl{(}-p-\tfrac{1}{2}\bigr{)}\,K_{p}(\omega)\,\omega^{1/2}\,\eta^{\,p+1/2}}{\sqrt{\pi}}.$
  • For $p>-1/2$:
    $f_{X|Y}(x\mid y)=f_{\mathrm{IG}}(x\,;\,\sqrt{\eta^{2}+2\eta y/\omega},\omega\eta+2y),$
    $f_{Y}(y)=\frac{1}{\mathcal{Z}_{1}}\,\frac{y^{p-1/2}}{\sqrt{\omega\eta+2y}}\,\exp\left\{-\omega\sqrt{1+2y/(\omega\eta)}\right\},$
    with
    $\mathcal{Z}_{1}=\frac{\sqrt{2}\,\Gamma\bigl{(}p+1/2)\,K_{p}(\omega)\,\eta^{p}}{\sqrt{\pi}}.$
The Gibbs sampler iteratively alternates between:
  1. 1.
    Sample $Y\mid X=x$:
    • For $p>-1/2$, $Y\mid X=x\sim\text{Gamma}(p+1/2,1/x)$.
    • For $p<-1/2$, $Y\mid X=x\sim\text{Gamma}(-(p+1/2),x)$.
  2. 2.
    Sample $X\mid Y=y$:
    • For $p<-1/2$, $X\mid Y=y\sim\mathrm{InvGauss}(\sqrt{\omega\eta/(\omega+2\eta y)},\omega\eta)$.
    • For $p>-1/2$, $X\mid Y=y\sim\mathrm{InvGauss}(\sqrt{\eta^{2}+2\eta y/\omega},\omega\eta+2y)$.
Proposition 4: If $X\sim\text{GIG}(p,\omega,\eta)$ and $p>1$, then:
$X=_{d}Y+E,$
where $Y$ and $E$ are independent random variables such that
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{Y}(y)=\tau\,f_{\text{GIG}}(y\,;\,p-2,\omega,\eta)+(1-\tau)\,f_{\text{GIG}}(y\,;\,p-1,\omega,\eta)$
and $E\sim\text{Exp}(\omega/(2\eta))$.
Proposition 5: Let $G_{p}(x)$ be the CDF of the $\mathrm{GIG}(p,\omega,\eta)$. For half-integer $p$ such that $|p|>1/2$, the following recurrence holds:
$G_{p}(x)=\tau G_{p-2}(x)+(1-\tau)G_{p-1}(x)-\exp\left(-\frac{\omega x}{2\eta}\right)I_{p}(x),$